Score: 0 of 3 pts 15 of 25 (24 complete) HW Score: 90%, 45 of 50 End of Chapter 2.3 Question Help The graph to the right shows how many pounds of apples and pounds of cherries you and your neighbor can each pick in one week. For example, if you devote all of your time to picking apples and none of your time to picking cherries, you can pick 12 pounds of apples. If you devote all of your time to picking cheries, you can pick 72 pounds. At the same time, if your neighbor devotes all of her time to picking apples, she can pick 32 pounds of apples. If she devotes all of her time to picking cherries, she can pick 32 pounds. PPF You Suppose initially that you (Y) are consuming 8 pounds of apples and 24 pounds of cherries and that your neighbor (N) is consuming 4 pounds of apples and 28 pounds of cherries, as indicated in the graph. Then, suppose you and your neighbor specialize by each only picking the good for which you have a comparative advantage and trade. In particular suppose you trade your neighbor half of your production for half of what your neighbor produces. In the table below, first fill in production when specializing. (Enter numeric responses using integers.) You Your Neighbor PPFNegh Apples Cherries Apples Cherries Consumption without trade Production with Trade 24 28 0 8 16 24 32 40 48 56 64 72 80 Apples (pounds)
Answer #1

Your opportunity cost of producing apples = 72/12= 6

Neighbors opportunity cost of producing apples = 32/32= 1

Thus you have a comparative advantage in cherries and neighbor has a comparative advantage in apples.

Production with trade

You- 0 apples, 72 cherries

Neighbor- 32 apples, 0 cherries